# Policy Gradient-Driven Noise Mask

### v2

@@ 3 Experiments
Implementation Details [^1] https://github.com/ogrenenmakine/Policy-Gradient-Driven-Noise-Mask [^2] https://huggingface.co/ogrenenmakine/Policy-Gradient-Noise-Mask The training pipeline is configured through a set of hyper-parameters. The main model to train is Resnet-50 and the policy network is always a leight-weight network such as Resnet-10t. The batch size is 32 for each 8xV100 GPU with effective batch size is 256. Each training phase takes 90 epochs with SGD optimizer with learning rate 0.1, momentum 0.9 and weight decay 1e-4. The step learning rate scheduler reduce by 1/10 in 30 epochs cycle. The Resnet-10t policy network starts with 0.01 learning rate and same momentum and weight decay and using cosine annealing learning rate scheduler. The initial values for $\alpha_{dataset}$ and $\beta_{dataset}$ are random and sampled from normal distribution.
@@ Appendix A RadImageNet: Artifacts and The Refinement
The RadImageNet database contains a diverse collection of ultrasound images, many of which exhibit various artifacts including text annotations, markers, and colorful segmentation masks. To address this issue, we developed an algorithm that automatically processes the standard RadImageNet distribution by implementing stratified splitting and removing ultrasound artifacts. This refined version of RadImageNet used in the paper, along with its corresponding code, is available on GitHub [^3] https://github.com/ogrenenmakine/Refined-RadImagenet.
@@ Appendix B Enhanced MedMNIST
MedMNIST is a comprehensive collection of standardized biomedical images designed for various analytical tasks in the medical field. This dataset has been expanded to include three new subsets, broadening the range of imaging modalities and classification challenges available to researchers. These additions complement the existing MedMNIST collections, offering a more diverse set of resources for developing and evaluating machine learning models across various medical imaging applications. The enhanced MedMNIST collection, including both existing and new datasets, is accessible on Hugging Face  [^4] https://huggingface.co/datasets/ogrenenmakine/Enhanced-MedMNIST

### v4

@@ 3 Experiments
Implementation Details [^1] https://github.com/convergedmachine/Policy-Gradient-Driven-Noise-Mask [^2] https://huggingface.co/convergedmachine/Policy-Gradient-Noise-Mask The training pipeline is configured through a set of hyper-parameters. The main model to train is Resnet-50 and the policy network is always a leight-weight network such as Resnet-10t. The batch size is 32 for each 8xV100 GPU with effective batch size is 256. Each training phase takes 90 epochs with SGD optimizer with learning rate 0.1, momentum 0.9 and weight decay 1e-4. The step learning rate scheduler reduce by 1/10 in 30 epochs cycle. The Resnet-10t policy network starts with 0.01 learning rate and same momentum and weight decay and using cosine annealing learning rate scheduler. The initial values for $\alpha_{dataset}$ and $\beta_{dataset}$ are random and sampled from normal distribution.
@@ Appendix A RadImageNet: Artifacts and The Refinement
RadImageNet encompasses an extensive collection of biomedical images across three primary imaging modalities: CT, MRI and US. Image statistics are provided in Table 5. The dataset is reorganized and divided into three stratified splits: training, validation, and testing (Figure 8). Below is a summary based on the provided histogram and data distribution: Training Data: 1,048,575 samples Validation Data: 50,182 samples Test Data: 100,364 samples. The dataset’s split into train, validation, and test sets ensures robust model development and evaluation. The histograms highlight potential class imbalances that may need addressing through techniques like resampling or weighting during model training.
Figure 8: Histograms illustrating the frequency distribution of samples across categories for the Train (1,048,575 samples), Validation (50,182 samples), and Test (100,364 samples) datasets. The charts highlight potential class imbalances that may impact model training and evaluation.
Anatomy # Classes Avg. Class Total #
A: CT (292,353 total images)
Lung 6 25,421 152,528
Abdomen/pelvis 28 4,994 139,825
B: Ultrasound (389,885 total images)
Thyroid 2 46,300 92,599
Abdomen/pelvis 13 22,868 297,286
C: MRI (672,675 total images)
Knee 18 9,975 179,555
Shoulder 14 3,743 52,407
Spine 9 7,964 71,674
Ankle/foot 25 7,264 181,603
Abdomen/pelvis 26 3,513 91,348
Brain 10 4,467 44,671
Hip 14 3,673 51,417
Table 5: Image statistics for CT, US, and MRI studies across various anatomical regions.
A.1 Artifact Refinement Process
The presence of artifacts (Figure 9) in ultrasound images necessitates a multi-step refinement process to enhance image quality: Region of Interest (ROI) Isolation: The algorithm processes specific regions within the images to identify and isolate areas affected by artifacts. Thresholding with Otsu’s Method: Applied to create binary masks that distinguish foreground from background elements. Morphological Operations: Utilized to remove minor artifacts and noise, ensuring the preservation of essential image features. Normalized Cross Correlation Coefficient (NCC): Employed to assess similarity to known artifact templates, aiding in the identification of contaminated regions. Adaptive Thresholding and Inpainting: Dynamically identifies artifact-affected areas and employs the Telea algorithm for inpainting, effectively restoring image integrity. Pixel-Level Outlier Detection: Provides an additional layer of quality assurance by ensuring the final images are free from residual artifacts.
@@ A.1 Artifact Refinement Process
A.2 Automated Conversion Tool for RadImageNet
In conjunction with this study, we introduce an automated tool designed to transform the existing RadImageNet dataset  [^3] https://github.com/convergedmachine/Refined-RadImagenet.  [^4] https://huggingface.co/convergedmachine/RadImagenet. into its refined and stratified version without requiring manual intervention. This tool uses predefined masks to remove artifacts such as text annotations, markers, and segmentation masks that are commonly found in ultrasound images.
By automating the refinement process, this tool significantly reduces the time and effort required to prepare dataset for pretraining purposes. It ensures that researchers can focus on developing fairly comparable diagnostic tools.
@@ Appendix B Enhanced MedMNIST
MedMNIST is a comprehensive collection of standardized biomedical images designed for various analytical tasks in the medical field. This dataset has been expanded to include three new subsets, broadening the range of imaging modalities and classification challenges available to researchers. These additions complement the existing MedMNIST collections, offering a more diverse set of resources for developing and evaluating machine learning models across various medical imaging applications (Table 6). The enhanced MedMNIST collection, including both existing and new datasets, is accessible on Hugging Face  [^5] https://huggingface.co/datasets/convergedmachine/Enhanced-MedMNIST